OpenStudy (anonymous):

The first four terms of a sequence are shown below: 8, 5, 2, -1 Which of the following functions best defines this sequence? f(1) = 8, f(n + 1) = f(n) + 3; for n 1 f(1) = 8, f(n + 1) = f(n) - 5; for n 1 f(1) = 8, f(n + 1) = f(n) + 5; for n 1 f(1) = 8, f(n + 1) = f(n) - 3; for n 1

OpenStudy (freckles):

first of all do you have a common difference?

OpenStudy (freckles):

5-8=? 2-5=? -1-2=?

OpenStudy (freckles):

nevermind didn't know the answer choices

OpenStudy (freckles):

but you could have set this up as a(n)=8+(-3)(n-1)

OpenStudy (freckles):

anyways...recursion

OpenStudy (freckles):

\[f(n+1)=f(n)+3\] replace n with 1

OpenStudy (freckles):

\[f(1+1)=f(1)+3\]

OpenStudy (freckles):

\[f(2)=f(1)+3\] we are given f(1)=8 so f(2)=?

OpenStudy (freckles):

we are trying the first choice to see what happens we are trying to see if it's terms match up with the terms given

OpenStudy (anonymous):

6

OpenStudy (freckles):

f(2)=f(1)+3 f(2)=8+3 f(2)=11 this first choice will not work because our second term was suppose to be 5 correct?

OpenStudy (freckles):

we know then that this f(1)+positive number will not work the numbers are suppose to decrease

OpenStudy (freckles):

check out one of the choices with a minus sign

OpenStudy (anonymous):

B and D

OpenStudy (freckles):

it is definitely one of those

OpenStudy (freckles):

the other two sequences will give us an increasing sequence

OpenStudy (freckles):

and we have a decreasing one

OpenStudy (freckles):

f(1)-5=? f(1)-3=?

OpenStudy (freckles):

which one will give us the correct second term

OpenStudy (anonymous):

f1 was 8 correct? so.. 8-5 = 3 8-3 = 5

OpenStudy (freckles):

yep

OpenStudy (freckles):

so is 3 our second or is 5?

OpenStudy (anonymous):

D

OpenStudy (freckles):

8, 5, 2, -1 so last option is the correct choice :) yep f(1)=8 f(2)=8-3=5 f(3)=5-3=2 f(4)=2-3=-1 and so on...

OpenStudy (anonymous):

I have another question

OpenStudy (freckles):

ok

OpenStudy (anonymous):

the slope for this

OpenStudy (freckles):

slope=rise/run right?

OpenStudy (freckles):

rise is positive if you rise up rise is negative if you rise down - run is positive if you go right run is negative if you go left

OpenStudy (anonymous):

yes

OpenStudy (freckles):

so let's count from Y to X

OpenStudy (freckles):

let's count up from Y

OpenStudy (freckles):

like one sec i will drae

OpenStudy (freckles):



OpenStudy (freckles):

 look at that vertical line segment I drew